ze:

9th Grade Math A team in science class placed a chalk mark on the side of a wheel and rolled the wheel in a straight line until the chalk mark returned to the same position. The team then measured the distance the wheel had rolled and found it to be 25 cm. To the nearest tenth, what is the area of the wheel? A. 99.5 cm2 B. 19.6 cm2 C. 198.9 cm2 D.49.7 cm2

wardchl000:

Sorry I am terrible at Math but....HHmm What grade are you in?

ze:

This is 9th grade math.

wardchl000:

Sorry I am in 7th grade

wardchl000:

Math I mean I uh had school problems...

Mercury:

The diameter is not 25. If the chalk mark on the wheel went around in a complete circle, the distance it traces is the circumference, not the diameter. Therefore: C = pi * d = 25 therefore, d = 25/pi divide d/2 to get the radius. then plug the radius into A = pi * r^2